OpenStudy (anonymous):

using multiplication (show work) -8 = a/4

OpenStudy (unklerhaukus):

\[-8={a \over 4}\]times both sides by 4 \[-8\times 4={a \over 4}\times 4\] \[-32=a\]

OpenStudy (anonymous):

-8 = a/4 -8 (4) = a -32 = a

OpenStudy (anonymous):

^ they're correct :)

OpenStudy (anonymous):

a = 4 * (-8) = -32